OpenStudy (anonymous):

medal for anyone who is right! please help!! what is the domain of the function y= sqrt 3x+3

OpenStudy (anonymous):

@Capnkookcx @lizz123 @LiliaRG

OpenStudy (anonymous):

26: simplify by rationalizing the denominator 4sqrt6/sqrt30 27:(2sqrt 5 +3sqrt7)^3

OpenStudy (e.mccormick):

Well, the domain is the input. So, can you think of any input to that function that would cause a problem?

OpenStudy (anonymous):

im not sure if its negative or posative

OpenStudy (e.mccormick):

And is that \(y= \sqrt {3x+3}\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it x>=-1?

OpenStudy (e.mccormick):

Yah, you are just doing the end result.

OpenStudy (anonymous):

ok so i was correct?

OpenStudy (e.mccormick):

Yes.

OpenStudy (anonymous):

awesome :) could you help me with the other 2 please :)

OpenStudy (e.mccormick):

Do you know what rationalizing means?

OpenStudy (anonymous):

kinda, not really, could you explain it to me?

OpenStudy (e.mccormick):

Well, when you have any sort of a fraction with variables, roots, etc. you try to get it into a form where all the odd stuff is on the top or on the bottom. With roots, they generally want all of it on the top. So you multiply by what is needed to move things to the top.

OpenStudy (e.mccormick):

\(\dfrac{\sqrt{a}}{\sqrt{b}}\) In that example, you would want to get the \(\sqrt{b}\) to the top. \(\dfrac{\sqrt{a}}{\sqrt{b}}\cdot \dfrac{\sqrt{b}}{\sqrt{b}}\) \(\dfrac{\sqrt{ab}}{\sqrt{bb}}\) \(\dfrac{\sqrt{ab}}{b}\)

OpenStudy (anonymous):

ok

OpenStudy (e.mccormick):

So you need to do something like that to the \(\dfrac{4\sqrt{6}}{\sqrt{30}}\)

OpenStudy (anonymous):

times sqrt30 to both sides?

OpenStudy (e.mccormick):

Yes. Then simplify it.

OpenStudy (anonymous):

so i have sqrt 4 sqrt 6 sqrt30/sqrt 30 sqrt 30

OpenStudy (e.mccormick):

\(\dfrac{4\sqrt{6}}{\sqrt{30}}\cdot \dfrac{\sqrt{30}}{\sqrt{30}}\) \(\dfrac{4\sqrt{6}\cdot\sqrt{30}}{\sqrt{30}\cdot\sqrt{30}}\) Yah, that. Then simplify it.

OpenStudy (anonymous):

im not sure how to smiplify it

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

Well, do you know how to multiply roots?

OpenStudy (anonymous):

well from 2sqrt6 *sqrt 30 is where im stuck

OpenStudy (anonymous):

is it 12sqrt5/30?

OpenStudy (anonymous):